Are the following lines parallel, perpendicular, or neither?
L1 through (–4, –7) and (1, 3)
L2 through (2, 6) and (4, 10)

Answer by Alan3354(69443)

Are the following lines parallel, perpendicular, or neither?
L1 through (–4, –7) and (1, 3)
L2 through (2, 6) and (4, 10)
-----------------------
Find the slopes of the 2 lines.
slope m = (y2-y1)/x2-x1)
For L1: m = (3 - (-7))/(1 - (-4))
m1 = 10/5 = 2
-----------
For L2: m = (10-6)/(4-2)
m2 = 4/2 = 2
The slopes are the same, so they're parallel.